Weekly Science Report 4: LMG02-05 Southern Ocean GLOBEC

Cruise days 19 to 25 August inclusive


BG-232-0 Costa/Burns/Crocker – Foraging Ecology of Crabeater seals

BG-234-0 Fraser – Winter Foraging Ecology of Adélie Penguins

BG-235-0 Fritsen – Sea Ice Microbial Communities

OG-241-0 Smith/Martinson/Perovich – Optical Environment of the Western Antarctic Peninsula Region

BG-244-0 Quetin/Ross – Winter Ecology of Larval Krill: Quantifying their interaction with the Pack Ice habitat



Cruise overview to date:


The week began with the LMG following the NBP toward Process Station #2.  Ice conditions and experience led the Captains and Ice Pilot Vladimir Repin of the NBP to reject several likely spots, but we finally found a suitable location a bit further offshore than originally planned.  At the same time, air temperatures dropped to around -25°C with wind chills of -50 to -60°C not uncommon.  A whale was sighted soon after arriving at station, now dubbed “Ice Station Minke.”  Unfortunately, after occupying this spot for only 26 hours we were forced to relocate after the floe rapidly broke apart.  All personnel and equipment were brought back safely on board.  The following morning we found another suitable floe nearby and proceeded to set up camp again.  An unusual ice formation reminded some people of the famous sculpture “The Thinker,” so we called this spot “Ice Station Rodin.”  The remainder of the week has proceeded relatively uneventfully, with scuba diving, establishment of ice/snow thickness transects, CTD casts, and drift net deployments occurring.  We've finally caught some krill so things are looking up.  The week ended with a sighting of an elephant seal (likely a sub-adult male) in the open water aft of the ship.  This is extremely unusual to see, especially in the winter.  We are continuing to drift with the pack and are now just north of the Antarctic Circle.


Daily Activities:


19 August 2002

The LMG followed the NBP last evening towards Survey Station #41, which was the target spot for Process Station #2.  However, ice conditions dictated otherwise so we backtracked to a place more likely to have sufficiently stable ice.  This, too, proved fruitless so we finished up science operations at this temporary site before moving to the east.  The weather has definitely changed and it now feels like Antarctica in the winter.  Air temperature dropped to -25°C with winds in the 30-40 knot range.  A CTD was done during the early morning hours.  The ice team sampled a nearby ice floe and the divers went into the water for about an hour in their continuing search for Krill.   In the Late afternoon the L.M. Gould continued looking for a good site for the next long term station being led by the N.B. Palmer.  We hope to be able to pick out a site early tomorrow.


20 August 2002

The LMG continued its convoy with the NBP throughout yesterday looking for a good location for Process Site #2.  No penguins were sighted during our transit, and no one was especially anxious to go outside as the temperatures dropped and the wind chill was in the -50 to -60°C range.  Several possible locations were abandoned due to the presence of icebergs and evidence of their breaking up the ice. This morning we finally found a suitable site that was away from icebergs and provided good access to a stable ice floe.  While we were setting up the station, a Minke whale surfaced immediately behind the ship.  Many of us ran down and saw it surface at least 3 times before it left.  Given our good fortune we named our new home for the next 7-9 days Ice Station Minke. 


21 August 2002

Antarctica, generally deemed the harshest of continents, demonstrated this reputation to us today.   As daylight broke in a cloudless sky, we initiated science by deploying a CTD beneath 4 inches of ice that formed last night.  Despite the 20-knot winds and temperatures ranging from -20 to -25°C (-50°C with wind chill!), excited scientists scattered out on the ice to build Ice Station Minke. The ice team set out to collect a series of ice cores to investigate changes in composition of the ice structure, algal biomass and microbial communities.   Samples were collected despite the repeated freezing of ice cores to the barrel shaft.  Additionally, the ice crew solicited help from the sealers, the penguin folk and the nutrient guy to shovel two quadrants created to monitor how varying light intensities affect algal and microbial productivity. Bundled divers preferred the relative warmth of -2°C seawater to the potentials of freezing air exposure.  After setting up transect lines under the ice, divers used suction tubes to collect samples of phytoplankton growing on the undersurface. With a solid day of science complete, folks gathered on the bridge to watch the sunset and as it turns out, to discover a large lead had opened near the Station.  The difficulty of studying a dynamic ice system is that it is…dynamic.  Consequently, we watched the harshest aspect of our day unfold.  The ice crumbled and cracked, dislodged the ship and consequently, Ice Station Minke was destroyed.  We collected our gear off the ice and made plans to find a new station at first light tomorrow.


22 August 2002

After the breakup of ice station Minke, we hove to for the night, then found a suitable floe in the morning and started over again.  We are currently on a piece of ice with large flat areas ~50 cm thick crossed by several small ridges.  It was a beautiful clear, though cold day. Most of the day was spent setting up our new ice station “Rodin”.  The krill team made diving holes in the ice and the ice team drilled an ice transect. We continue to experience extremely cold but sunny weather.  The temperature remained around -20-25°C with wind chill down to - 50°C.  However, the cold did not prevent a friendly game of soccer on the ice. In the evening we were treated to a spectacular sunset with the red sky growing more and more brilliant for several hours after the sun dipped below the horizon.


23 August 2002

The LMG continues to occupy the new Process Station #2, conducting a full suite of operations including: ice coring, snow/ice thickness transects, scuba diving, and CTD casts.  We are still without a penguin sighting for the past few days, but seals are occasionally seen.  Today it's getting cloudier, and therefore warmer as the downward longwave radiation increases. After 3 days of sunny, but cold weather, we finally got a break in the activities so we could take a cruise group photo.  The ice dynamics team deployed a simple ice buoy and the ice biota team collected ice cores. In the evening the krill team deployed nets from the ships stern to fish for krill.


24 August 2002

We continue to make excellent use of Process Station #2, conducting a full suite of operations including: ice coring, snow/ice thickness transects, scuba diving, and CTD casts.  Three scuba dives were accomplished and the second of three ice buoys were deployed yesterday.  We are still seeing rather sparse fauna in this location with no seals, whales, or birds seen and only a few krill.  Occasionally a ctenophore will float by during a CTD cast.  We are hopeful of maintaining this floe, but the forecast is for winds in the next day or two. The day started with clear sunny skies, but became increasingly overcast as the day continued.


25 August 2002

We continue to occupy Process station #2 (Rodin).  The weather has changed considerably with 15-20 knot winds coming from the W to NW.  It is warmer with temperatures around -9°C.  We've finally caught some krill using drift nets deployed through the dive holes and left out overnight.  The Ice team is re-measuring their ice/snow thickness transect line and taking optics measurements.  Blustery weather has prevented the seal team from working on the lone crabeater hauled out in our vicinity.  Several other seals have been sighted in the water nearby, though.  We have still not seen any penguins lately. Late in the evening while conducting the evening CTD, a sub-adult male southern elephant surfaced in the open water aft of the ship. It was definitely a sub-adult male elephant seal, as the marine mammal group, several of whom work with elephant seals, positively identified it. A sighting of an elephant seal in the pack ice this far south is extremely unusual, especially in the winter.


Individual group reports: 


BG-232 – Seal Foraging Ecology- (Costa, Shaffer, Barnes, McDonald, and Kuhn)

Due to the cold weather very few if any seals have been sighted hauled out and none were sighted at the Ice Station.  We have taken the time to analyze samples collected during our intense period of tag deployments.  This includes completing blood hemoglobin, blood volume and blood count analysis. So far the animals tagged on this cruise have remained in the general area where they were tagged.  However, the leopard seal appears to have a more restricted movement pattern than the crabeater seals.  One of the seals tagged at the first ice station has apparently moved into the southern area of Crystal Sound. Unfortunately, 3 of the 11 tags that were deployed are currently not transmitting.  This could be because the animals are in the water, or due to tag malfunction.  We are hopeful that with the warming trend the seals will be more likely to haul out increasing the probability of getting data from these tags.  We are still planning to capture seals for stomach sampling as well as to collect additional body composition data.  In the meantime we are analyzing our samples as well as helping the other science groups with ice coring, buoy deployments and diver tending.


BG-234 – Penguin Foraging Ecology (Heidi Geisz, Brett Pickering)

Penguins remained scarce and very few birds were seen at all during the past week.  While transiting to Ice Station Minke a single Adélie Penguin was spotted on a flow near the LMG’s path through the ice.  However, by the time the ship was in position to deploy personnel over the side a thick bank of fog rolled in and the bird was visually lost.  Also at approximately the same time 3 to 5 Adélies were noted swimming in a lead, but were not spotted leaving the water.  Four Emperor penguins were observed swimming in small leads the morning before we moved in to Ice Station Minke.  While at our two ice stations this week the only birds spotted were several Snow Petrels and a single Antarctic Petrel.  The penguin team remained quite busy, however, assisting BG 235 with their ice coring work and BG 244 with dive operations.


BG-235 - Sea Ice Microbial Communities (Adkins, Blees, Cunningham, Hartsough, Fritsen)

CTD operations. We (in conjunction with Raytheon and ECO personnel) have been running morning (09:00) and evening (21:00) CTD’s to a depth of 300 meters to track the changes in the water column properties that affect and are affected by ice cover dynamics.   These casts are accomplished by running the Seabird ensemble package (CTD, fluorometer, beam c, oxygen) out of the Baltic room to the stern where a hole is being maintained by the ship's prop wash.  While at this second process station we have recorded the deep mixed layers ranging from 75 to 98 meters.  Several casts, however, have exhibited a weak T-S step (temperature increase of 0.03°C, salinity increase of 0.013 psu) well above the winter mixed layer at 18 to 30 meters.  Before we set into the ice we did manage to conduct one CTD cast with the rosette.  This cast showed chlorophyll a averaging 0.03 micrograms per liter.


Ice Transects: In conjunction with OG-241 (Elder and Claffey) and with the much appreciated help from members of the seal and penguin projects (Geisz, Pickering and Shaffer) we drilled ice thickness transects (n=130) across ridged areas and relatively flat areas of the floe that we are parked next to.  Portions of these transects were chosen to coincide with the Krill program (BG-244 Quetin) in order to facilitate comparisons of ice properties and ablation dynamics that can determined from working from the surface and those that are only obtainable from diving operations.


Coring: We have been sampling ice cores every day since our arrival at this station.  Thus, far we have sampled 26 cores and processed 25 meters of ice for nutrients, biota and productivity assays. Ice cores have ranged from 35 cm to over 4 meters in length.    In addition we have been processing samples from stations past.   Average chlorophyll content in ice cores collected using the “toss the stick in the wind and core where it lands technique” (a procedure employed to help ensure a nested random sampling regime) continue to range between 2 to 50 micrograms per liter.  Of the cores taken, approximately 30% have visible algal coloration somewhere in the ice profile.   The vertical location of the algal “bands” has varied with these colorations occurring mostly toward the bottom half of the ice cores.   Bacterial production and photosynthesis experiments continue in an effort to determine the timing and controls on production and biomass accumulation in the varied ice habitats that are sometimes undergoing major temperature and salinity and irradiance shifts as the early spring season begins.


OG-241 Optical Environment of the Western Antarctic Peninsula Region (Elder and Claffey)

In the last week we have been positioned at 2 different stations. Ice Station “Minke” and “Rodin”. The first station was short lived as it broke up approximately 26 hours after it was founded. On this station we performed 70 meters of ice transect and collocated ice cores at 2 sites along the transects. This was all done with BG235 and BG244.


Ice station Rodin has thus far stayed together. We conducted 130meters of transect line to measure ice thickness and snow depth (assistance from the penguin and seal teams was greatly appreciated). We hope to re-measure this transect once or twice before leaving to measure ice growth as the temperatures have been below -20°C for a majority of the time on station. The ice is thickening which is evident by the skeletal/growth layer on the bottom of the ice cores.  These are new crystals that grow down off the bottom of the ice sheet.


Ice coring has been more sporadic on this station than others as we have made an attempt to get all the ice types in the area, not necessarily along one of our transect lines.  We sampled the top 4 meters of some older/thicker ice that had 110 cm of snow on top of it.  The snow contained many icy crust layers that indicate either rain or melt from times much warmer than the current temperatures. We also have sampled in the rubbled ice between the floes in search of the brown algae layers on the bottom sides that we see while breaking ice. These seem to be in the rubble/blocks between the floes and not on the flat ice that comprises the floes.


An ice-drifting buoy was deployed upon one of the thicker ice floes. This buoy transmits back hourly position, air temperature and barometric pressure.  The NBP spotted the buoy that was deployed last week. It had drifted 38nm NNE from where we had left it 5 days earlier.  This correlates with our drift while on station as up until Saturday evening, we also drifted NNE.


While on station Passive microwave measurements have been recorded on the flat ice sheet that we are parked in which is approximately 40cm thick.  We began the task of our ice structure work in the freezer van in the hold of the ship. Although it is warmer inside the freezer than the outside temperatures, we all agree we would rather be outside.


BG-244 - Winter Ecology of Larval Krill: Quantifying their interaction with the Pack Ice habitat (Quetin, Hessell, Willis, Boch, Oakes, Johnston and Dovel)

This week has been varied and productive.  Tuesday we dove from the stern of the ship and found enough krill for length and stage frequency measurements.  These krill were much younger and smaller than our previous catches further south, with few F6 and many F4 and F5 stages.  After the dive we moved to a more suitable location to set up Ice Station Minke.  On Wednesday we secured an underwater transect line and collected 6 samples of the under-ice community from different ice surfaces.  Unfortunately, by 19:30 LT Ice Station Minke had sounded, and all aboard were impressed by how quickly the ice surrounding the floe split away from it.  Later, we were able to retrieve our Scott Tent that was left on the ice when we pulled away.


Ice Station Rodin was established on Thursday.  Set up Thursday morning went quickly, and we were able to set up 2 underwater transect lines in the afternoon.  We are presently established at Ice Station Rodin and have drifted generally in a northeasterly direction.  Krill larvae feeding directly on under-ice surfaces continue to be scarce.  We have spent much of our dive time collecting under-ice surface samples to assess that environment as a feeding area for larval krill in comparison to the water column.  As the station has drifted northward we have seen an increase in macro zooplankton during our dives, and yesterday some schools of older krill moved through the area.  These schools were exciting to watch but difficult to photograph.  One of our divers did manage to collect enough of the larger krill for an Instantaneous Growth Rate experiment.  In addition, Saturday night we used Ice Station Rodin to tow two 1-m nets.  Sunday morning we were rewarded with a large number of AC0 and AC1 krill.  Ice Station Rodin has been our most productive station to date, and we hope the good weather continues for the next several days.


We would like to thank all aboard who have helped us sustain our effort this week, especially given the temperatures of -24°C ambient (-55°C with wind chill) that we encountered.  Dealing with the “freezing factor” required extra people, and we could have not accomplished what we did without your help.  Thanks!



Submitted by:

Daniel Costa, Chief Scientist and

Karl Newyear, Marine Projects Coordinator

LMG 02-05